# Learning the Boundary of Solvability: Aligning LLMs to Detect Unsolvable Problems

### v2

@@ Abstract
Ensuring large language model (LLM) reliability requires distinguishing objective unsolvability (inherent contradictions) from subjective capability limitations (tasks exceeding model competence). Current LLMs often conflate these dimensions, leading to hallucinations in which they return confident answers to inherently unsolvable queries. To address this issue, we propose a multi-domain dataset containing both solvable and unsolvable questions, UnsolvableQA, together with an alignment framework, UnsolvableRL. First, we construct UnsolvableQA by “Reverse Construction” that systematically injects logical contradictions into otherwise valid reasoning chains. Second, we introduce UnsolvableRL, a reinforcement learning paradigm that balances objective unsolvability detection with calibrated confidence under capability limits. Empirically, our approach achieves near-perfect unsolvability detection ($90\%$ detection rate) and boosts solvable reasoning accuracy from 43.4% to 69.4% on Qwen3-4B-Instruct. Crucially, we identify a data–training interaction: strict alignment constraints induce Capability Collapse without unsolvable data, but act as a regularizer for rigor when such data are included, thereby improving overall robustness. Our code and data are available at https://github.com/sfasfaffa/unsolvableQA.
Learning the Boundary of Solvability: Aligning LLMs to Detect Unsolvable Problems
@@ 1 Introduction
As Large Language Models (LLMs) achieve strong performance on complex reasoning tasks (chen2025towards; wei2022chain; guo2025deepseek; jaech2024openai), a key research direction (qin2025largelanguagemodelsmeet) is ensuring they recognize their own limitations and behave reliably. Earlier work on confidence calibration (xiong2023can; li2024confidence; yoon2025reasoning; fisch2022calibrated) during reasoning underscores the need for models to know when not to answer. Further, studies on LLMs’ capability boundaries (chen2024unlocking) examine the self-awareness of their limitations to avoid unproductive reasoning on problems they are likely to fail (zhang2025self; mei2025reasoning; kalai2025languagemodelshallucinate; chen2025aware). More recently, TruthRL (wei2025truthrl) has encouraged LLMs to refuse when uncertain through a ternary reward mechanism.
@@
First, to support rigorous analysis of unsolvability, we develop a fully automated synthesis pipeline and introduce UnsolvableQA, a dataset of paired solvable and unsolvable instances spanning puzzles and mathematics. For programmatically verifiable puzzle tasks, we construct provably unsolvable instances by injecting code-level logical contradictions. For mathematics problems without reliable programmatic verification, inspired by lightman2024lets; ling2023deductive, we propose a “Reverse Construction” method: starting from a solvable problem and its correct rationale, we introduce a controlled contradiction into the rationale and then prompt LLMs to generate an unsolvable problem that matches the resulting flawed reasoning.
Then, to train models that remain robust on both solvable and unsolvable inputs, we propose UnsolvableRL, a reinforcement-learning (RL) method based on GRPO (shao2024deepseekmath). Unlike prior approaches that conflate refusal types, we define 3 distinct RL objectives: (1) Inherent unsolvability rewards generating an explicit refusal tag when solvability is detected; (2) False unsolvability Penalty discourages erroneous refusals by penalizing labeling solvable problems as unsolvable, thereby requiring explicit solvability checks before refusal; (3) Capability calibration uses a mechanism that rewards refusal only when the problem difficulty exceeds the LLMs’ current capability.
Empirical results show a synergistic effect: UnsolvableRL elevates the unsolvable problem detection rate of Qwen3-4B from 40.3% to 90.7% (near-perfect) and boosts its solvable problem accuracy from 43.4% to 69.4%, achieving an overall score of 80.1%. Our ablation study reveals three key insights: (1) models cannot learn to detect unsolvable problems without explicit unsolvable training data; (2) penalizing false refusals without such data causes “capability collapse”, where models stop refusing altogether; (3) when paired with unsolvable data, the penalty acts as a regularizer, trading a small drop in detection for stronger reasoning accuracy, a Pareto improvement over baselines.
@@ 2 Related Work
As the reliance of Large Language Models on data quality increases, Automated Dataset Construction (ADC) has emerged as a prominent research direction (liu2024automatic; wang2023data). In the realm of reasoning, recent efforts on automatic data construction primarily focus on ensuring data solvability and verifiability through formal or programmatic means. Representative works such as Enigmata (chen2025enigmata) and BIGGSM (chen2024unlocking) have significantly enriched the resources for solvable reasoning tasks by employing synthetic puzzles and difficulty-graded annotations, respectively. Concurrently, research on verifying the correctness of rationale traces has also been advancing (ling2023deductive; chen2025towards).
In contrast, progress in dataset construction regarding problem unsolvability has been relatively slow. Earlier, SQuAD 2.0 (rajpurkar2018knowdontknowunanswerable) examined whether questions are answerable given a long context, but it does not adequately evaluate refusal behavior in complex logical-reasoning settings. Consequently, for reasoning tasks, existing data-synthesis efforts focus mainly on solvable instances while largely neglecting unsolvable questions, leaving a significant research gap.
Figure 2: The UnsolvableQA data collection pipeline. (a) Mathematical Data Collection: We employ a “Reverse Construction” method where contradictions are injected into the solution path of solvable problems. The resulting instances are verified by an LLM to ensure they lead to a contradiction (unsolvable) rather than a valid solution. (b) Puzzle Data Collection: We utilize domain-specific programmatic generators and automated deterministic verifiers (e.g., SAT solvers, DFS) to rigorously classify randomly generated instances into solvable and unsolvable sets.
@@
Beyond dataset construction, Reinforcement Learning (RL) (ouyang2022traininglanguagemodelsfollow; schulman2017proximalpolicyoptimizationalgorithms; rafailov2024directpreferenceoptimizationlanguage) has become a central approach for aligning LLMs with desired behaviors such as safety and truthfulness. Recent RL algorithms for reasoning, such as GRPO (shao2024deepseekmath) and DAPO (yu2025dapoopensourcellmreinforcement), further enhance models’ reasoning abilities (sheng2025hybridflow). Concurrent work also leverages RL to improve reasoning efficiency under cost constraints (tu2025learning; team2025kimi; guo2025deepseek; wang2025aspoasymmetricimportancesampling; chen2025aware). In parallel, advances such as TruthRL (wei2025truthrl) introduce objectives that encourage models to refuse likely hallucinations, and research on self-awareness enables models to decline answering when appropriate (zhang2025self; mei2025reasoning; kalai2025languagemodelshallucinate; Steyvers2025).
@@ 3.1 Mathematics Problem Generation via Reverse Construction
Validation Process. Based on this plan, LLMs then reverse-engineers a new problem statement that preserves the stylistic and complexity characteristics of the seed problem while embedding the designed contradiction. Finally, we implement a two-tier verification protocol: the generated problem is first presented to DeepSeek-R1 to autonomously detect contradictions; if model fails to independently identify the contradiction, we proceed to the second tier. In this stage, we explicitly provide the model with both the final unsolvable problem statement and the predefined contradiction plan, thereby supplementing it with the intended contradiction information beyond the problem text alone. The model is then prompted to reassess the problem’s solvability given this complete contextual guidance. Only instances that are ultimately verified as unsolvable through this two-tier procedure are retained.
@@ 3.2 Logic Puzzle Generation
Game24. Solvability is checked by exhaustively enumerating all binary expression trees with exact rational arithmetic (Python Fraction) to avoid floating-point error. Unsolvable instances are obtained by rejection sampling, retaining number sets for which the exhaustive search proves that no solution exists.
Hamiltonian Cycle/Path. We encode graph traversability as a Boolean Satisfiability problem and verify randomly generated graphs using SAT solvers (MiniSat backend via pysat). Unsolvable instances are constructed by injecting structural impossibilities, such as removing critical edges to create bottlenecks or disconnecting components.
Hitori. We formulate Hitori as a constraint satisfaction problem and, using python-constraint to generate puzzles with unique solutions. For unsolvable instances, we inject conflicting constraints, for example forcing a row or column to contain duplicate numbers that cannot be removed without violating adjacency rules.
Maze Navigation. Base mazes are generated via randomized Depth-First Search (DFS) to ensure nontrivial structure. To obtain unsolvable instances, we analyze all valid paths in a solvable maze and place an obstacle at a critical junction shared by every path, then verify unsolvability using Breadth-First Search (BFS).
@@ 3.3 Dataset Statistics and Quality Control
We construct UnsolvableQA to maintain approximately balanced instances per domain (Table 1). The dataset comprises 899 test problems (458 solvable, 441 unsolvable) and 637 training problems (348 solvable, 289 unsolvable). While training prompts follow templated formats and pool Maze difficulties, test prompts are diversified and preserve fine-grained splits (including distinct AIME years [^1] AIME year ranges: train 1983–2023; test 2024–2025.). All instances undergo rigorous quality control: logic puzzles are verified via SAT/CSP/graph solvers, mazes via deterministic search, and math problems through multi-pass LLM verification. We also introduce two OOD testsets in Appendix A.2.
@@ 4.1 Problem Formulation
Specifically, to satisfy three distinct behavioral objectives, our training framework optimizes a policy $\pi_{\theta}$ as follows. First, for solvable and easy instances within the model’s capabilities ($x\in\mathcal{S}\cap\mathcal{E}$), the objective is Reasoning Accuracy, requiring the model to produce correct reasoning chains and final answers. Second, for inherently unsolvable instances ($x\in\mathcal{U}\cap\mathcal{E}$), the objective is Unsolvability Detection: the model must explicitly recognize the inconsistency and output a dedicated tag (e.g., <unsolvable>) instead of hallucinating a solution, which we treat as an objective correctness task. Third, for hard yet solvable instances ($x\in\mathcal{H}$) that exceed the model’s effective reasoning capacity (as indicated by low internal confidence), the objective is Capability Calibration, where the model should prudently abstain (e.g., output <beyond_capacity>) to reduce false positives.
@@ 4.2 Group-Relative Policy Optimization
where $\mu_{\mathbf{R}}$ and $\sigma_{\mathbf{R}}$ are the mean and standard deviation of the rewards $\mathbf{R}=\{R(y_{j}\mid x)\}_{j=1}^{G}$ within the group. This group-based normalization can stabilize training due to reward-scale differences across solvable and unsolvable domains.
@@ 4.3 Reward Design
Dataset Metric Deepseek GPT-5.1 Gemini-3 Instruct + UnsolRL Instruct + UnsolRL
@@
Mean 98.0 90.0 93.5 40.0 39.5 38.9 77.9
AIME 24-25 S / U 85.0 / 63.8 61.7 / 48.9 95.0 / 34.1 38.3 / 33.5 35.4 / 52.7 67.9 / 25.3 69.6 / 62.2
Mean 74.4 55.3 64.5 35.9 44.1 46.6 65.9
Overall S / U 88.4 / 87.7 45.9 / 67.6 97.4 / 86.0 23.0 / 43.4 25.5 / 79.8 43.4 / 40.3 69.4 / 90.7
Mean 88.1 56.7 91.7 33.2 52.7 41.9 80.1
Table 2: Detailed performance comparison. For each dataset, the first row reports Solvable Accuracy (S) / Unsolvable Detection Rate (U), and the second row reports their Mean (M). We bold the global best score across all models for each metric. Qwen3-4B + UnsolRL achieves SOTA performance in Unsolvable Detection (Overall U) and Game24 Mean accuracy.
@@
It evaluates the correctness of LLMs’ judgments on verifiably solvable instances. (1) A Correct Response (+1+1) is assigned when the model outputs the correct answer for a solvable instance. (2) For Incorrect Reasoning (0) on solvable instances, such as producing an incorrect final answer despite performing the calculation, the model receives a neutral reward. This asymmetric scheme emphasizes reasoning capability by rewarding successful problem solving while not explicitly penalizing unsuccessful but genuine reasoning attempts on solvable tasks.
@@
This component evaluates how accurately the model identifies inherently contradictory or ill-posed instances. (1) A Correct-Detect Response (+1+1) is assigned when the model explicitly outputs the canonical tag unsolvable for such an instance. (2) To prevent degenerate behavior, we impose a False Unsolvability Penalty ($\rho=-0.5$): if the model incorrectly declares a solvable problem unsolvable, it receives a negative reward. As derived in our decision-theoretic analysis in Appendix A.3, this penalty is mathematically critical to prevent the model from collapsing into a local optimum of universal rejection. It effectively expands the decision threshold, encouraging the model to attempt reasoning even under moderate uncertainty. (3) For Error-Detect Reasoning (0) on unsolvable instances, such as hallucinating a solution instead of outputting unsolvable, the model receives a neutral reward.

### v3

# Learning the Boundary of Solvability: Aligning Large Language Models to Detect Unsolvable Problems

@@ Abstract
Ensuring large language model (LLM) reliability requires distinguishing objective unsolvability (inherent contradictions) from subjective capability limitations (tasks exceeding model competence). Current LLMs often conflate these dimensions, leading to hallucinations in which they return confident answers to inherently unsolvable queries. To address this issue, we propose a multi-domain dataset containing both solvable and unsolvable questions, UnsolvableQA, together with an alignment framework, UnsolvableRL. First, we construct UnsolvableQA by “Reverse Construction” that systematically injects logical contradictions into otherwise valid reasoning chains. Second, we introduce UnsolvableRL, a reinforcement learning paradigm that balances objective unsolvability detection with calibrated confidence under capability limits. Empirically, our approach achieves robust unsolvability detection ($85\%$ detection rate) and boosts solvable reasoning accuracy from 43.4% to 69.4% on Qwen3-4B-Instruct. Crucially, we identify a data–training interaction: strict alignment constraints induce Capability Collapse without unsolvable data, but act as a regularizer for rigor when such data are included, thereby improving overall robustness. Our code and data are available at https://github.com/sfasfaffa/unsolvableQA.
Learning the Boundary of Solvability:
Aligning Large Language Models to Detect Unsolvable Problems
@@ 1 Introduction
As Large Language Models (LLMs) achieve strong performance on complex reasoning tasks (Chen et al., 2025c; Wei et al., 2022; Guo et al., 2025; Jaech et al., 2024; Hu et al., 2024), a key research direction (Qin et al., 2025) is ensuring they recognize their own limitations and behave reliably. Earlier work on confidence calibration (Xiong et al., 2023; Li et al., 2024; Yoon et al., 2025; Fisch et al., 2022) during reasoning underscores the need for models to know when not to answer. Further, studies on LLMs’ capability boundaries (Chen et al., 2024) examine the self-awareness of their limitations to avoid unproductive reasoning on problems they are likely to fail (Zhang et al., 2025; Mei et al., 2025; Kalai et al., 2025; Chen et al., 2025b). More recently, TruthRL (Wei et al., 2025) has encouraged LLMs to refuse when uncertain through a ternary reward mechanism.
@@
First, to support rigorous analysis of unsolvability, we develop a fully automated synthesis pipeline and introduce UnsolvableQA, a dataset of paired solvable and unsolvable instances spanning puzzles and mathematics. For mathematics, inspired by Lightman et al. (2024); Ling et al. (2023), we propose a “Reverse Construction” method that injects controlled contradictions into valid rationales to generate matching unsolvable problems. For programmatically verifiable puzzles, we construct instances via code-level logical contradictions. Notably, our logic puzzle pipeline enables the rapid generation of infinite instances with arbitrary difficulty, and we will fully release our construction code upon acceptance.
Then, to train models that remain robust on both solvable and unsolvable inputs, we propose UnsolvableRL, a reinforcement-learning (RL) method based on GRPO (Shao et al., 2024). Unlike prior approaches that conflate refusal types, we define 3 distinct RL objectives: (1) Inherent unsolvability rewards generating an explicit refusal tag when solvability is detected; (2) False unsolvability Penalty discourages erroneous refusals by penalizing labeling solvable problems as unsolvable, thereby requiring explicit solvability checks before refusal; (3) Capability calibration uses a mechanism that rewards refusal only when the problem difficulty exceeds the LLMs’ current capability.
Empirical results show a synergistic effect: UnsolvableRL elevates the unsolvable problem detection rate of Qwen3-4B from 38.8% to 87.5% and boosts its solvable problem accuracy from 43.4% to 69.4%, achieving an overall score of 78.6%. Our ablation study reveals three key insights: (1) models cannot learn to detect unsolvable problems without explicit unsolvable training data; (2) penalizing false refusals without such data causes “capability collapse”, where models stop refusing altogether; (3) when paired with unsolvable data, the penalty acts as a regularizer, trading a small drop in detection for stronger reasoning accuracy, a Pareto improvement over baselines.
@@ 2 Related Work
As the reliance of Large Language Models on data quality increases, Automated Dataset Construction (ADC) has emerged as a prominent research direction (Liu et al., 2024; Wang et al., 2023). In the realm of reasoning, recent efforts on automatic data construction primarily focus on ensuring data solvability and verifiability through formal or programmatic means. Representative works such as Enigmata (Chen et al., 2025a) and BIGGSM (Chen et al., 2024) have significantly enriched the resources for solvable reasoning tasks by employing synthetic puzzles and difficulty-graded annotations, respectively. Concurrently, research on verifying the correctness of rationale traces has also been advancing (Ling et al., 2023; Chen et al., 2025c).
In contrast, progress in dataset construction regarding problem unsolvability has been relatively slow. Earlier, SQuAD 2.0 (Rajpurkar et al., 2018) examined whether questions are answerable given a long context, but it does not adequately evaluate refusal behavior in complex logical-reasoning settings. Consequently, for reasoning tasks, existing data-synthesis efforts focus mainly on solvable instances while largely neglecting unsolvable questions, leaving a significant research gap.
Figure 2: The UnsolvableQA data collection pipeline. (a) Mathematical Data Collection: It employs a “Reverse Construction” method where contradictions are injected into the solution path of solvable problems. The resulting instances are verified by an LLM to ensure they lead to a contradiction (unsolvable) rather than a valid solution. (b) Puzzle Data Collection: It utilizes domain-specific programmatic generators and automated deterministic verifiers (e.g., SAT solvers, DFS) to rigorously classify randomly generated instances into solvable and unsolvable sets.
@@
Beyond dataset construction, Reinforcement Learning (RL) (Ouyang et al., 2022; Schulman et al., 2017; Rafailov et al., 2024) has become a central approach for aligning LLMs with desired behaviors such as safety and truthfulness. Recent RL algorithms for reasoning, such as GRPO (Shao et al., 2024) and DAPO (Yu et al., 2025), further enhance models’ reasoning abilities (Sheng et al., 2025). Concurrent work also leverages RL to improve reasoning efficiency under cost constraints (Tu et al., 2025; Team et al., 2025; Guo et al., 2025; Wang et al., 2025; Chen et al., 2025b). In parallel, advances such as TruthRL (Wei et al., 2025) introduce objectives that encourage models to refuse likely hallucinations, and research on self-awareness enables models to decline answering when appropriate (Zhang et al., 2025; Mei et al., 2025; Kalai et al., 2025; Steyvers et al., 2025).
@@ 3.1 Mathematics Problem Generation via Reverse Construction
Validation Process. Based on this plan, LLMs then reverse-engineers a new problem statement that preserves the stylistic and complexity characteristics of the seed problem while embedding the designed contradiction. Finally, we implement a two-tier verification protocol: the generated problem is first presented to DeepSeek-R1 to autonomously detect contradictions; if model fails to independently identify the contradiction, we proceed to the second tier. In this stage, we explicitly provide the model with both the final unsolvable problem statement and the predefined contradiction plan, thereby supplementing it with the intended contradiction information beyond the problem text alone. The model is then prompted to reassess the problem’s solvability given this complete contextual guidance. Only instances verified as unsolvable are retained, and we manually verified the final test set.
@@ 3.2 Logic Puzzle Generation
Game24.
Solvability is checked by exhaustively enumerating all binary expression trees with exact rational arithmetic (Python Fraction) to avoid floating-point error. Unsolvable instances are obtained by rejection sampling, retaining number sets for which the exhaustive search proves that no solution exists.
Hamiltonian Cycle/Path.
We encode graph traversability as a Boolean Satisfiability problem and verify randomly generated graphs using SAT solvers (MiniSat backend via pysat). Unsolvable instances are constructed by injecting structural impossibilities, such as removing critical edges to create bottlenecks or disconnecting components.
Hitori.
We formulate Hitori as a constraint satisfaction problem and, using python-constraint to generate puzzles with unique solutions. For unsolvable instances, we inject conflicting constraints, for example forcing a row or column to contain duplicate numbers that cannot be removed without violating adjacency rules.
Maze Navigation.
Base mazes are generated via randomized Depth-First Search (DFS) to ensure nontrivial structure. To obtain unsolvable instances, we analyze all valid paths in a solvable maze and place an obstacle at a critical junction shared by every path, then verify unsolvability using Breadth-First Search (BFS).
@@ 3.3 Dataset Statistics and Quality Control
We construct UnsolvableQA to maintain approximately balanced instances per domain (Table 1). The dataset comprises 899 test problems (458 solvable, 441 unsolvable) and 637 training problems (348 solvable, 289 unsolvable). While training prompts follow templated formats and pool Maze difficulties, test prompts are diversified and preserve fine-grained splits (including distinct AIME years [^1] AIME year ranges: train 1983–2023; test 2024–2025.). All instances undergo rigorous quality control: logic puzzles are verified via SAT/CSP/graph solvers, mazes via deterministic search, and math problems through multi-pass LLM verification and manual checks. We also introduce two OOD testsets in Appendix A.2.
@@ 4.1 Problem Formulation
Specifically, to satisfy three distinct behavioral objectives, our training framework optimizes a policy $\pi_{\theta}$ as follows. First, for solvable and subjectively easy instances within the model’s capabilities ($x\in\mathcal{S}\cap\mathcal{E}$), the objective is Reasoning Accuracy, requiring the model to produce correct reasoning chains and final answers. Second, for inherently unsolvable instances ($x\in\mathcal{U}\cap\mathcal{E}$), the objective is Unsolvability Detection: the model must explicitly recognize the inconsistency and output a dedicated tag (e.g., <unsolvable>) instead of hallucinating a solution, which we treat as an objective correctness task. Third, for subjectively hard yet solvable instances ($x\in\mathcal{H}$) that exceed the model’s effective reasoning capacity (as indicated by low internal confidence), the objective is Capability Calibration, where the model should prudently abstain (e.g., output <beyond_capacity>) to reduce false positives.
@@ 4.2 Group-Relative Policy Optimization
where $\mu_{\mathbf{R}}$ and $\sigma_{\mathbf{R}}$ are the mean and standard deviation of the rewards $\mathbf{R}=\{R(y_{j}\mid x)\}_{j=1}^{G}$ within the group. This group-based reward facilitates efficient and stable training, helps effectively distinguish different responses to the same problem, and eliminates the need for critic and reward models that occupy massive GPU memory.
@@ 4.3 Reward Design
Dataset Metric Deepseek-V3.2-R GPT-5.1-Low Gemini-3 Instruct + UnsolvableRL Instruct + UnsolvableRL
@@
Mean 98.0 90.0 93.5 40.0 39.5 (-0.5) 38.9 77.9 (+39.0)
AIME 24-25 S / U 85.0 /40.3 61.7 / 42.5 95.0 / 21.2 38.3 / 21.2 35.4 / 28.7 67.9 / 14.8 69.6 / 40.4
Mean 62.7 52.1 58.1 29.8 32.1 (+2.3) 41.4 55.0 (+13.6)
Overall S / U 88.4 / 84.4 45.9 / 66.6 97.4 / 84.1 23.0 / 41.7 25.5 / 76.4 43.4 / 38.8 69.4 / 87.5
Mean 86.1 56.2 90.8 32.4 50.9(+18.5) 41.1 78.6 (+37.5)
Table 2: Detailed performance comparison. For each dataset, the first row reports Solvable Accuracy (S) / Unsolvable Detection Rate (U), and the second row reports their Mean (M). Values in parentheses indicate the absolute improvement over the base Instruct model. We bold the global best score across all models. Qwen3-4B + UnsolvableRL achieves SOTA performance in Unsolvable Detection (Overall U) and Game24 Mean accuracy.
@@
It evaluates the correctness of LLMs’ judgments on verifiably solvable instances. (1) A Correct Response (+1+1) is assigned when the model outputs the correct answer for a solvable instance. (2) For Incorrect Reasoning (0) on solvable instances, such as producing an incorrect final answer despite performing the calculation, the model receives a neutral reward (0). This ensures that honest calculation errors are not penalized, maintaining the model’s incentive to attempt reasoning.
@@
This component evaluates how accurately the model identifies inherently contradictory or ill-posed instances. (1) A Correct-Detect Response (+1+1) is assigned when the model explicitly outputs the canonical tag unsolvable for such an instance. (2) To prevent degenerate behavior, we impose a False Unsolvability Penalty ($\rho=-0.5$): if the model incorrectly declares a solvable problem unsolvable, it receives a negative reward. This penalty effectively expands the decision threshold, encouraging the model to attempt reasoning even under moderate uncertainty. As derived in our decision-theoretic analysis in Appendix A.5, this term is mathematically critical to prevent the model from collapsing into a local optimum of universal rejection. Moreover, our empirical ablation on $\rho$ selection identifies $\rho=-0.5$ as the optimal value; while stricter penalties yield marginal gains in solvable accuracy, they significantly impair unsolvable detection capabilities, making this value the most effective trade-off. We refer to this penalty mechanism as $P$ in our subsequent experimental analysis and tables. (3) For Error-Detect Reasoning (0) on unsolvable instances, such as hallucinating a solution instead of outputting unsolvable, the model receives a neutral reward.
@@
Figure 4: Detailed Ablation Metrics. Breakdown of Solvable Accuracy (S, Red), Unsolvable Detection Rate (U, Blue), and Mean Performance (M, Green) across different configurations. The w/o UnsData (Unsolvable Data) w/ P setting shows a catastrophic drop in U for the 1.7B model (4.2%), indicating capability collapse. Notably, w/ UnsData w/o P achieves the highest Mean (M) score by maximizing U. The full method (w/ UnsData w/ P) demonstrates a valuable trade-off, boosting S while maintaining a high U, optimizing for practical scenarios.